# Regression-Based Proximal Causal Inference

### v1

@@ Abstract
In observational studies, identification of causal effects is threatened by the potential for unmeasured confounding. Negative controls have become widely used to evaluate the presence of potential unmeasured confounding thus enhancing credibility of reported causal effect estimates. Going beyond simply testing for residual confounding, proximal causal inference (PCI) was recently developed to debias causal effect estimates subject to confounding by hidden factors, by leveraging a pair of negative control variables, also known as treatment and outcome confounding proxies. While formal statistical inference has been developed for PCI, these methods can be challenging to implement in practice as they involve solving complex integral equations that are typically ill-posed. In this paper, we develop a regression-based PCI approach, employing a two-stage regression via familiar generalized linear models to implement the PCI framework, which completely obviates the need to solve difficult integral equations. In the first stage, one fits a generalized linear model (GLM) for the outcome confounding proxy in terms of the treatment confounding proxy and the primary treatment. In the second stage, one fits a GLM for the primary outcome in terms of the primary treatment, using the predicted value of the first-stage regression model as a regressor which as we establish accounts for any residual confounding for which the proxies are relevant. The proposed approach has merit in that (i) it is applicable to continuous, count, and binary outcomes cases, making it relevant to a wide range of real-world applications, and (ii) it is easy to implement using off-the-shelf software for GLMs. We establish the statistical properties of regression-based PCI and illustrate their performance in both synthetic and real-world empirical applications.
@@ 1 Introduction
Identification of causal effects can be challenging in observational studies because the standard exchangeability or treatment unconfoundedness conditions upon which causal inference commonly relies is threatened by the presence of confounding by hidden factors [1]. To achieve unconfoundedness in practice typically requires a complete understanding of key causes of the treatment selection mechanism that also predict the outcome in view, and that such common causes of the treatment and outcome variables can be measured with sufficient accuracy that one can fully account for them empirically. However, even in carefully designed observational studies, where all sources of confounding are known, key drivers of confounding can rarely if ever be measured without error. As a result, the most one can hope for in practice is that measured confounders are (error-prone) proxies of underlying confounders, thus residual unmeasured confounding may be present even in the most well-designed observational studies. For instance, in studies of the causal impact of HIV treatment, pre-treatment CD4 count measurement is commonly used to control for confounding by indication as it is invariably used by physicians for treatment initiation decision-making [2, 3]. However, CD4 count measurements are notoriously error-prone and are at most imperfect snapshots of the evolving immune system status driving disease progression. As such, CD4 count measurement is a proxy of immune function, the underlying source of confounding which remains unmeasured[4]. Pre-exposure outcome measurements sometimes provide good candidate proxies of confounding, e.g. Asthma rates at a county level measured the week before PM2.5 measurements are good confounding proxies for the causal effect of PM2.5 exposure on the county’s Asthma rate in the subsequent week in a typical air pollution time series study [5].
@@
In the continuous outcomes case, under linear models for NCO $W$ and primary outcome $Y$, Tchetgen Tchetgen et al. (2023) establish that the corresponding proximal g-computation algorithm can be implemented by following a two-stage least squares procedure [13]. A key limitation of the proposed two-stage least squares approach is that it does not easily extend to binary, polytomous or count outcomes for which linear models are generally inappropriate.
In this paper, we extend proximal g-computation for a point treatment via a two-stage regression approach which accommodates continuous, count and binary $W$ and $Y$ via generalized linear models (GLMs) with canonical link (identity, log and logit) functions that appropriately account for the outcome type. The paper is organized as follows. In the next section, we briefly review causal Directed Acyclic Graphs (DAGs) encoding key assumptions of the proximal causal learning framework and we review the two-stage least square estimation given in Tchetgen Tchetgen et al. for continuous outcome and outcome proxy variables. Next, we consider more general settings, where the primary outcome or the outcome proxy is either a count, binary or polytomous variable and we propose appropriate two-stage regression-based PCI approaches to identify and estimate the corresponding causal effect. For each combination of link functions, we provide a detailed formulation of the corresponding estimation procedure, and formally establish its validity. Detailed proofs and derivations are relegated to the Appendix. For illustration, we re-analyze data from the Study to Understand Prognoses and Preferences for Outcomes and Risks of Treatments (SUPPORT)[17], an observational study that evaluates the causal effect of right heart catheterization (RHC) on survival, length of hospitalization, intensity of care, and costs.
@@ 2.1 Notation and Causal Structure
Assumption (2) encodes the condition that as valid treatment confounding proxy, $Z$ must not associated with $Y$ given $(A,X,U)$; while Assumption (3) encodes the condition that as valid outcome confounding proxy $W$ must not be associated with $(A,Z)$ given $(X,U)$. Of note, $Z$ and $A$ may be causally associated and so can $W$ and $Y$ provided (2) and (3) hold. Notably, proxies that are conditionally independent of both $A$ and $Y$ given $X,U$ may be specified as either treatment or outcome confounding proxy.
In the SUPPORT study, the treatment $A$ indicates receiving RHC treatment, and the outcome $Y$ is survival time up to 30 days which we will in turn consider as a continuous variable and its dichotomized version indicating 30-day survival. The study also collected rich information encoded in 73 covariates including basic demographics (e.g. age, sex, race, education, income and insurance status), estimated probability of survival, comorbidity, vital signs, physiological status and functional status. Ten covariates measuring a patient’s overall physiological status were measured from a blood test drawn within 24 hours in the ICU: serum sodium, serum potassium, serum creatinine, bilirubin, albumin, PaO2/(.01*FiO2) ratio, PaCO2, serum PH (arterial), white blood cell count, and hematocrit. These variables may be subject to substantial measurement error and as single snapshots of the underlying physiological state over time can be considered potential confounding proxies.
Following Tchetgen Tchetgen et al. (2023), we specify biomarker measurements (pafi1, paco21) and (ph1, hema1) as proxies $Z$ and $W$ respectively, and include all remaining physiological status measurements, medical history, and baseline variables such as age and gender in $X$.
In subsequent sections, consistent with DAG (2) we suppress all measured covariates $X$ in order to simplify the exposition of the proposed two-stage regression approach. We focus on model specification, fitting and estimation of causal effects via two-stage regression with GLMs, deferring the details of large sample behavior and statistical inference to Section 2.6. Note, all proposed results accommodate DAG (1), which represents the special case where $W$ and $Y$ are conditionally independent given $U$ and so are $A$ and $Z$. In Section 2.6, we also adapt the approach to the more general setting depicted in DAG (3) by incorporating measured covariates $X$.
@@ 2.2 Continuous $(Y,W)$
We briefly review the two-stage least-squares approach for continuous $(Y,W)$ in which case the identity link function is appropriately specified [13]. Specifically, suppose that in our analysis of the SUPPORT data, $W$ denotes continuous measurements (ph1, hema1) and $Y$ denotes the length of survival up to 30 days, treated as a continuous variable. In this case, it may be reasonable to assume that the following linear structural models hold:
@@
The second equation thus establishes that $E[W|A,Z]$ and $E[Y|A,Z]$ are both linearly related with $E[U|A,Z]$, implying equation (8) below, which identifies the causal effect $\beta_{a}$ upon rearranging a few terms as previously reported in Tchetgen Tchetgen et al.:
@@
The condition $\alpha_{u}\neq 0$ encodes the requirement that $W$ is relevant for, and therefore associated with $U$. Result 1 states that if $E[W|A,Z]$ were known, one could in principle estimate $\beta_{a}$ by regressing $Y$ on $(A,E[W|A,Z])$ via standard least-squares fitting. In practice, as $E[W|A,Z]$ will not generally be known, Tchetgen Tchetgen et al. proposed to estimate it in the first stage by fitting a regression of $W$ on $(A,Z)$, to obtain fitted values for each observation, $\hat{E}[W|A,Z]$ from, say regression equation (9) below, although more flexible models can in principle be used (see Section 2.6 for further details):
@@
As established in Tchetgen Tchetgen et al. the resulting estimator $\hat{\beta}_{a}$ is consistent and asymptotically normal under the stated assumptions, provided that $W$ and $Z$ are $U-$relevant such that the coefficient $\alpha_{z}^{*}\neq 0$ also. Furthermore, as they point out, this two-stage least-squares procedure is reminiscent of an analogous two-stage least-squares approach routinely used in the instrumental variable literature. This analogy is useful as it implies that one can implement the above proximal two-stage linear regression approach and obtain corresponding inference using off-the-shelf software such as the R function ”ivreg”[13]. Moreover, it is important to note that the regression equations (4) and (5) are not contingent upon the presence of direct causal links from $Z$ to $A$ and from $W$ to $Y$. This indicates that the same regression procedure to estimate $\beta_{a}$ applies to DAG (1). Having reviewed two-stage least-squares, we are now ready to consider settings where $Y$ and $W$ may no longer be continuous, in which case linear models may no longer be appropriate. While we focus the exposition on the canonical case where both variables $W$ and $Y$ are of the same type (e.g. both are counts or binary), we later discuss how the methods can be adapted to accommodate settings where they are of different types.
@@ 2.3 Count $(Y,W)$
$\beta_{a}=\mathrm{log}\left(\frac{E[Y^{a=1}|U]}{E[Y^{a=0}|U]}\right)=\mathrm{log}\left(\frac{E[Y|A=1,Z,U]}{E[Y|A=0,Z,U]}\right)$
@@
We generally therefore expect that under correct model specification, $\hat{\beta}_{a}$ will be consistent for $\beta_{a}$ and approximately normal in large samples. See Section 2.6 for further details on obtaining inference about $\beta_{a}$. This result formally establishes proximal two-stage Poisson regression as a practical approach for implementing PCI in the context of count primary outcome and outcome confounding proxy variables. Next, we consider the case of binary outcomes.
@@ 2.4 Binary $(Y,W)$
To motivate the setting, revisiting the SUPPORT study, suppose that only a dichotomized (at the median) version of $W$ is made available to the analyst, and that the primary outcome $Y$ is likewise a binary indicator of 30-day survival. In such a case, one may wish to posit models that appropriately account for the natural scale of the outcomes. As the $\mathrm{logit}$ link is routinely used for binary outcomes, thus suppose that a logistic structural equation model is appropriate to model such data. In this setting, we show that some care is required in specifying an appropriate two-stage regression approach which accounts for the nonlinear logit link function.
@@
$\beta_{a}$ $=\mathrm{log}\left(\frac{\mathrm{Pr}(Y^{a=1}=1|W,U)/\mathrm{Pr}(Y^{a=1}=0|W,U)}{\mathrm{Pr}(Y^{a=0}=1|W,U)/\mathrm{Pr}(Y^{a=0}=0|W,U)}\right)$
$=\mathrm{logit}(\mathrm{Pr}(Y=1|A=1,Z,W,U))-\mathrm{logit}(\mathrm{Pr}(Y=1|A=0,Z,W,U))$
@@
Result 3 states that $\beta_{a}$ can be estimated by fitting a logistic regression of $A$ and $W$ adjusting for $\mathrm{logit}({\mathrm{Pr}}(W=1|A,Z,Y=0))$ as a covariate. As the latter is not known, as before one may proceed in two stages, first by estimating $\mathrm{logit}\widehat{\mathrm{Pr}}(W=1|A,Z,Y=1))$ from a logistic regression for the conditional probability of $W=1$ conditional on $(A,X,Y)$. The two-stage regression procedure to estimate $\beta_{a}$ is thus as follows:
@@
    Perform the second-stage logistic regression $Y\sim A+\mathrm{logit}\widehat{\mathrm{Pr}}(W=1|A,Z,Y=1))+W$ to obtain $\hat{\beta}_{a}$.
The above model specification excludes interactions between $(Y,W)$ and $(A,Z)$, which restricts specifying more flexible regression functions. However, such a restriction can be relaxed by allowing for certain interactions as detailed in Section A.3.2 of the Appendix. We should also note that the approach requires no modification even if $W$ is a priori known not to cause $Y$ as in DAG (1).
@@ 2.5 Polytomous $(Y,W)$
The proposed approach readily extends with a slight modification to accommodate polytomous outcome and outcome proxy by fitting polytomous logistic regression models rather than simple logistic regression models. Let us consider a scenario where $W$ has $K+1$ categories $(0,\ldots,K)$ and $Y$ has $(T+1)$ categories $(0,\ldots,T)$. We set 00 as the reference level for $W$ and $Y$ and proceed with the following parameterization:
@@
$U=E[U|A,Z,W=0,Y=0]+\epsilon;\ E[\epsilon]=0;\ \epsilon\perp\!\!\!\perp(A,Z)|(Y=0,W=0);$ (20)
the distribution of $\epsilon|(Y=0,W=0)$ is unrestricted.
@@
$=\tilde{\beta}_{0t}+\beta_{at}^{*}A+\beta_{ut}^{*}\sum_{k=1}^{K}\mathrm{logit}(\mathrm{Pr}(W=k|A,Z,Y=1))+\sum_{k=1}^{K}\tilde{\beta}_{wtk}I(W=k)\ \textrm{, where $\beta_{at}^{*}=\beta_{at}$.}$ (21)
@@
    Compute $\sum_{k=1}^{K}\mathrm{logit}(\widehat{\mathrm{Pr}}(W=k|A,Z,Y=1))=\sum_{k=1}^{K}\left(\hat{\alpha}_{0k}^{*}+\hat{\alpha}_{ak}^{*}A+\hat{\alpha}_{zk}^{*}Z+\sum_{t=1}^{T}\hat{\alpha}_{ykt}^{*}\right)$ using the estimated coefficients from the first-stage regression.
@@
    Perform the second-stage polytomous logistic regression $Y\sim A+\sum_{k=1}^{K}\mathrm{logit}\widehat{\mathrm{Pr}}(W=k|A,Z,Y=1))+W$ to obtain $\hat{\beta}_{at}$.
2.6 Unification of Proximal inference for GLMs
W Data Type Y Data Type
Continuous (Identity Link)
Count (Log Link)
Binary (Logit Link)
Continuous (Identity Link)
Linear $W\sim A+Z$
$S=E[W|A,Z]$
Linear $Y\sim A+S$
Linear $W\sim A+Z$
$S=E[W|A,Z]$
Poisson $Y\sim A+S$
Linear $W\sim A+Z+Y$
$S=E[W|A,Z,Y=1]$
Logistic $Y\sim A+S$
Count (Log Link)
Poisson $W\sim A+Z$
$S=\mathrm{log}(E[W|A,Z])$
Linear $Y\sim A+S$
Poisson $W\sim A+Z$
$S=\mathrm{log}(E[W|A,Z])$
Poisson $Y\sim A+S$
Poisson $W\sim A+Z+Y$
$S=\mathrm{log}(E[W|A,Z,Y=1])$
Logistic $Y\sim A+S$
Binary (Logit Link)
Logistic $W\sim A+Z$
$S=\mathrm{logit}(\mathrm{Pr}(W=1|A,Z))$
Linear $Y\sim A+S+W$
Logistic $W\sim A+Z$
$S=\mathrm{logit}(\mathrm{Pr}(W=1|A,Z))$
Poisson $Y\sim A+S+W$
Logistic $W\sim A+Z+Y$
$S=\mathrm{logit}(\mathrm{Pr}(W=1|A,Z,Y=1))$
Logistic $Y\sim A+S+W$
Table 1: Estimation procedures for each combination of three link functions. $S$ denotes proximal control variable for $U$.
In this Section, we provide a unified formulation of the proximal two-stage regression approach under different combinations of GLMs for $Y$ and $W$, respectively.
Importantly, each instance delivers a procedure that produces a consistent estimator of the causal effect of interest, whether or not there is a direct causal link from either $Z$ to $A$ or from $W$ to $Y$. Further details are provided in the Appendix.
For simplicity, thus far the exposition has ignored measured confounders $X$ and interaction terms, we now describe the ease with which they can be incorporated in the proposed methodology, by simply including them as covariates in regressions at both stages. Specifically, suppose that $Y$ and $W$ are both continuous and therefore modeled using linear models compatible with DAG (3) which now also involves measured confounders $X$, thus the first stage linear model may take the form:
while the second stage may take the form:
A similar approach applies to all other cases under different combinations of link functions, including those involving logit links where $(Y,W)-(A,Z,X)$ interactions might be considered (see Appendix for more details).
We evaluated the finite sample performance of the proximal two-stage regression approach in extensive simulation studies focused primarily on the binary settings, which we report in Section A.7 of the Appendix. Our study verifies that in the logit link case, the estimated parameter $\hat{\beta}_{a}$ obtained with two-stage regression converges to the true parameter $\beta_{a}$ under correct specification of the underlying structural models $(Y,W)$. In contrast, standard logistic regression adjusting for $X$ only is severely biased in its inferences about ${\beta}_{a}$. For inference, the bootstrap may be convenient in practice; however, we also provide analytic expressions for the asymptotic variance of parameter estimates obtained via proximal two-stage regression which we have implemented in R package ”pci2s”. The proof of asymptotic normality and standard error computation are also provided in Section A.8 of the Appendix.

### v3

@@ Abstract
Negative controls are increasingly used to evaluate the presence of potential unmeasured confounding in observational studies. Beyond the use of negative controls to detect the presence of residual confounding, proximal causal inference (PCI) was recently proposed to de-bias confounded causal effect estimates, by leveraging a pair of treatment and outcome negative control or confounding proxy variables. While formal methods for statistical inference have been developed for PCI, these methods can be challenging to implement as they involve solving complex integral equations that are typically ill-posed. We develop a regression-based PCI approach, employing two-stage generalized linear regression models (GLMs) to implement PCI, which obviates the need to solve difficult integral equations. The proposed approach has merit in that (i) it is applicable to continuous, count, and binary outcomes cases, making it relevant to a wide range of real-world applications, and (ii) it is easy to implement using off-the-shelf software for GLMs. We establish the statistical properties of regression-based PCI and illustrate their performance in both synthetic and real-world empirical applications.
@@ 1 Introduction
Identification of causal effects can be challenging in observational studies because the standard exchangeability or treatment unconfoundedness conditions upon which causal inference commonly relies is threatened by the presence of confounding by hidden factors [1]. To achieve unconfoundedness in practice typically requires a complete understanding of key causes of the treatment selection mechanism that also predict the outcome in view, and that such common causes of the treatment and outcome variables can be measured with sufficient accuracy that one can fully account for them empirically. However, even in carefully designed observational studies, where all sources of confounding are known, key drivers of confounding can rarely if ever be measured without error. As a result, the most one can hope for in practice is that measured confounders are (error-prone) proxies of underlying confounders, thus residual unmeasured confounding may be present even in the most well-designed observational studies. For instance, in studies of the causal impact of HIV treatment, pre-treatment CD4 count measurement is commonly used to control for confounding by indication as it is invariably used by physicians for treatment initiation decision-making [2, 3]. However, CD4 count measurements are notoriously error-prone and are at most imperfect snapshots of the evolving immune system status driving disease progression. As such, in scenarios where CD4 count directly affects treatment, CD4 count measurement is a proxy of the underlying source of confounding which remains unmeasured [4]. Pre-exposure outcome measurements sometimes provide good candidate proxies of confounding, e.g. Asthma rates at a county level measured the week before PM2.5 measurements are good confounding proxies for the causal effect of PM2.5 exposure on the county’s Asthma rate in the subsequent week in a typical air pollution time series study [5].
@@
In the continuous outcomes case, under linear models for NCO $W$ and primary outcome $Y$, Tchetgen Tchetgen et al. (2023) establish that the corresponding proximal g-computation algorithm can be implemented by following a two-stage least-squares procedure [13]. A key limitation of the proposed two-stage least-squares approach is that it does not easily extend to binary, polytomous or count outcomes for which linear models are generally inappropriate.
In this paper, we extend proximal g-computation for a point treatment via a two-stage regression approach which accommodates continuous, count and binary $W$ and $Y$ via generalized linear models (GLMs) with canonical link function (identity, log and logit) that appropriately account for the outcome type. The paper is organized as follows. In the next section, we briefly review causal Directed Acyclic Graphs (DAGs) encoding key assumptions of the proximal causal learning framework and we review the two-stage least-squares estimation given in Tchetgen Tchetgen et al. (2023) for continuous outcome and outcome proxy variables. Next, we consider more general settings, where the primary outcome or the outcome proxy is either a count, binary or polytomous variable and we propose appropriate two-stage regression-based PCI approaches to identify and estimate the corresponding causal effect. For each combination of link functions, we provide a detailed formulation of the corresponding estimation procedure, and formally establish its validity. Detailed proofs and derivations are relegated to the Appendix. We also provide annotated R and SAS codes to execute the two-stage regression procedure proposed in this paper in the Appendix A.10. Additionally, we have developed R package pci2s, to conveniently estimate causal effects with consistent standard error estimates and corresponding confidence intervals. For illustration, we re-analyze data from the Study to Understand Prognoses and Preferences for Outcomes and Risks of Treatments (SUPPORT)[17], an observational study that evaluates the causal effect of right heart catheterization (RHC) on survival, length of hospitalization, intensity of care, and costs.
@@ 2.1 Notation and Causal Structure
Assumption (2) encodes the condition that as valid treatment confounding proxy, $Z$ must not be associated with $Y$ given $(A,X,U)$; while Assumption (3) encodes the condition that as valid outcome confounding proxy $W$ must not be associated with $(A,Z)$ given $(X,U)$. Of note, $Z$ and $A$ may be causally associated and so can $W$ and $Y$ provided (2) and (3) hold. Notably, proxies that are conditionally independent of both $A$ and $Y$ given $X,U$ may be specified as either treatment or outcome confounding proxy.
In subsequent sections, consistent with DAG (2) we suppress all measured covariates $X$ in order to simplify the exposition of the proposed two-stage regression approach. We focus on model specification, fitting and estimation of causal effects via two-stage regression with GLMs on outcomes $Y$ and $W$ with identical link functions, deferring details on the use of different link functions, large sample behavior, statistical inference and simulation studies to Section A.1 of the Appendix. The approach readily extends to the setting depicted in DAG (3) by incorporating measured covariates $X$ in both regression models of the two-stage approach. All results hold in DAG (1) a submodel of DAG (3), where $W$ and $Y$, and $A$ and $Z$ are conditionally independent given $U$, respectively. For inference, one may either use the nonparametric bootstrap, or an empirical estimate of the asymptotic variance we derive in Section A.9 of the Appendix.
@@ 2.2 Continuous $(Y,W)$
We briefly review the two-stage least-squares approach for continuous $(Y,W)$ in which case the identity link function is appropriately specified [13]. It may be reasonable to assume that the following linear structural models hold:
@@
These equations thus establish that $E[W|A,Z]$ and $E[Y|A,Z]$ are both linearly related with $E[U|A,Z]$, implying equation (8) below, which identifies the causal effect $\beta_{a}$ upon rearranging a few terms as previously reported in Tchetgen Tchetgen et al.:
@@
The condition $\alpha_{u}\neq 0$ encodes the requirement that $W$ is relevant for, and therefore associated with $U$. Result 1 states that if $E[W|A,Z]$ were known, one could in principle estimate $\beta_{a}$ by regressing $Y$ on $(A,E[W|A,Z])$ via standard least-squares fitting. In practice, as $E[W|A,Z]$ will not generally be known, Tchetgen Tchetgen et al. proposed to estimate it in the first stage by fitting a regression of $W$ on $(A,Z)$, to obtain fitted values for each observation, $\hat{E}[W|A,Z]$ from, say regression equation (9) below, although more flexible models can in principle be used (see Section A.1 for further details):
@@
As they establish, the resulting estimator $\hat{\beta}_{a}$ is consistent and asymptotically normal under the stated assumptions, provided that $W$ and $Z$ are $U-$relevant such that the coefficient $\alpha_{z}^{*}\neq 0$. Furthermore, as they point out, this two-stage least-squares procedure is reminiscent of an analogous two-stage least-squares approach routinely used in the instrumental variable literature. This analogy is useful as it implies that one can implement the above proximal two-stage linear regression approach and obtain corresponding inference using off-the-shelf software such as the R function “ivreg”[13]. Moreover, it is important to note that the regression equations (4) and (5) are not contingent upon the presence of direct causal links from $Z$ to $A$ and from $W$ to $Y$. This indicates that the same regression procedure to estimate $\beta_{a}$ applies to DAG (1).
Regarding the violation of key assumptions for valid proxies, identification of the treatment effect would not be possible without invoking additional assumptions. For example, if exclusion condition 3 $W\perp\!\!\!\perp(A,Z)\mid X,U$ is violated and $A$ has a direct causal effect on $W$, the first-stage linear regression equation may be
$E[W|A,Z]=\alpha_{0}+\alpha_{a}A+\alpha_{u}E[U|A,Z]$ (10)
instead of equation 7. And if the main effect of $A$ in $E[U|A,Z]$ is linear, for example $E[U|A,Z]=\gamma_{0}+\gamma_{a}A+\gamma_{u}U$, it may take the form
$E[W|A,Z]=(\alpha_{0}+\alpha_{u}\gamma_{0})+(\alpha_{a}+\alpha_{u}\gamma_{a})A+\alpha_{u}\gamma_{z}Z.$ (11)
In this scenario, the result of second-stage linear regression equation 6 of $Y$ on $A$ and $E[W|A,Z]$ from 10 will not be able to distinguish $\beta_{a}^{*}$ from $\beta_{a}$ and $\beta_{a}-\beta_{u}\frac{\alpha_{a}}{\alpha_{u}}$.
Having reviewed two-stage least-squares, we are now ready to consider settings where $Y$ and $W$ may no longer be continuous, in which case linear models may no longer be appropriate. While we focus the exposition on the canonical case where both variables $W$ and $Y$ are of the same type (e.g. both are counts or binary), we later discuss how the methods can be adapted to accommodate settings where they are of different types.
@@ 2.3 Count $(Y,W)$
$\mathrm{log}\left(\frac{E[Y^{a=1}|U]}{E[Y^{a=0}|U]}\right)=\mathrm{log}\left(\frac{E[Y|A=1,Z,U]}{E[Y|A=0,Z,U]}\right)=\beta_{a}$
@@
This result formally establishes proximal two-stage Poisson regression as a practical approach for implementing PCI in the context of count primary outcome and outcome confounding proxy variables. For Poisson regression with overdispersion, one may either use a standard sandwich estimator for robust variance estimation or opt for a Negative Binomial regression. Both methods have been implemented in the R package “pci2s”. Next, we consider the case of binary outcomes.
@@ 2.4 Binary $(Y,W)$
Suppose $W$ and $y$ are dichotomous. In such a case, one may wish to posit models that appropriately account for the natural scale of the outcomes. As the $\mathrm{logit}$ link is routinely used for binary outcomes, thus suppose that a logistic structural equation model is appropriate to model such data. In this setting, we show that some care is required in specifying an appropriate two-stage regression approach which accounts for the nonlinear logit link function.
@@
$\mathrm{log}\left(\frac{\mathrm{Pr}(Y^{a=1}=1|W,U)/\mathrm{Pr}(Y^{a=1}=0|W,U)}{\mathrm{Pr}(Y^{a=0}=1|W,U)/\mathrm{Pr}(Y^{a=0}=0|W,U)}\right)$
$=\mathrm{logit}(\mathrm{Pr}(Y=1|A=1,Z,W,U))-\mathrm{logit}(\mathrm{Pr}(Y=1|A=0,Z,W,U))=\beta_{a}$
@@
Result 3 states that $\beta_{a}$ can be estimated by fitting a logistic regression of $A$ and $W$ adjusting for $\mathrm{logit}({\mathrm{Pr}}(W=1|A,Z,Y=1))$ as a covariate. As the latter is not known, one may proceed in two stages as before , first by estimating $\mathrm{logit}(\widehat{\mathrm{Pr}}(W=1|A,Z,Y=1))$ from a logistic regression for the conditional probability of $W=1$ conditional on $(A,X,Y)$. The two-stage regression procedure to estimate $\beta_{a}$ is thus as follows:
@@
    Perform the second-stage logistic regression $Y\sim A+\mathrm{logit}(\widehat{\mathrm{Pr}}(W=1|A,Z,Y=1))+W$ to obtain $\hat{\beta}_{a}$.
The above model specification excludes interactions between $(Y,W)-U$, which restricts specifying more flexible regression functions. However, such a restriction can be relaxed by allowing for certain interactions as detailed in Section A.4.2 of the Appendix. We should also note that the approach requires no modification even if $W$ is a priori known not to cause $Y$ as in DAG (1).
@@ 2.5 Polytomous $(Y,W)$
The proposed approach readily extends with a slight modification to accommodate polytomous outcome and outcome proxy by fitting polytomous logistic regression models rather than simple logistic regression models. Let us consider a scenario where $W$ has $K+1$ categories $(0,\ldots,K)$, $Y$ has $T+1$ categories $(0,\ldots,T)$ and scalar $U$. We set 00 as the reference level for $W$ and $Y$ and proceed with the following parameterization:
@@
$U|A,Z,Y=0,W=0\sim E[U|A,Z,Y=0,W=0]+\epsilon;\ E[\epsilon]=0;$ (22)
$\epsilon\perp\!\!\!\perp(A,Z)|Y=0,W=0;\textrm{ the distribution of $\epsilon|Y=0,W=0$ is unrestricted.}$.
@@
$=\beta_{0t}^{*}+\beta_{at}^{*}A+\beta_{ut}^{*}\sum_{k=1}^{K}\mathrm{logit}(\mathrm{Pr}(W=k|A,Z,Y=0))+\sum_{k=1}^{K}\tilde{\beta}_{wtk}I(W=k)\ \textrm{, where $\beta_{at}^{*}=\beta_{at}$.}$ (23)
@@
    Compute $\sum_{k=1}^{K}\mathrm{logit}(\widehat{\mathrm{Pr}}(W=k|A,Z,Y=0))=\sum_{k=1}^{K}\left(\hat{\alpha}_{0k}^{*}+\hat{\alpha}_{ak}^{*}A+\hat{\alpha}_{zk}^{*}Z\right)$ using the estimated coefficients from the first-stage regression.
@@
    Perform the second-stage polytomous logistic regression $Y\sim A+\sum_{k=1}^{K}\mathrm{logit}(\widehat{\mathrm{Pr}}(W=k|A,Z,Y=0))+W$ to obtain $\hat{\beta}_{at}$.
Multidimensional $U$ can be accommodated with certain rank conditions satisfied for high-dimensional $(Z,W)$. In Section A.4.3 of the Appendix, we establish a more general version of Result 4 and outline the corresponding regression procedure for multidimensional $U$.